OpenStudy (anonymous):

Write the equation in slope intercept form. The line with slope -1 that passes through (3,5)?

OpenStudy (anonymous):

use https://www.desmos.com/calculator

pooja195 (pooja195):

y=mx+b we know that the slope is -1 y=-x+b now plug in the given coordinates to find the b part of the equation 5=-(3)+b solve for b

OpenStudy (anonymous):

y = -x + 8 (-x is equivalent to -1)

OpenStudy (anonymous):

y = mx + b slope = -1 coordinates are (3,5) 3 will be plugged in the "x" because that is the x-coordinate and 5 will be plugged in the "y" because y is the y-coordinate. 3 will become -3 since the line has a slope of -1 which will be m. 5 = -(3) + b +3 = +3 + b (you would add both sides by 3 since -(3) would make it -3. That cancels out the three on the right side and you get your answer on the left) 8 = b